Exercise 7-4 Basic Segmented Income Statement [LO7-4]

Royal Lawncare Company produces and sells two packaged products—Weedban and Greengrow. Revenue and cost information relating to the products follow:

Product

Weedban Greengrow
Selling price per unit $ 11.00 $ 31.00
Variable expenses per unit $ 2.80 $ 15.00
Traceable fixed expenses per year $ 132,000 $ 40,000

Common fixed expenses in the company total $100,000 annually. Last year the company produced and sold 40,000 units of Weedban and 23,500 units of Greengrow.

Required:

Prepare a contribution format income statement segmented by product lines.
Answer #1
Total Weedban Greengrow
Sales 1,168,500 (40,000*11) = 440,000 (23,500*31) = 728,500
Variable costs 464,500 (40,000*2.80) = 112,000 (23,500*15) = 352,500
Contribution margin 704,000 328,000 376,000
Traceable fixed expenses 172,000 132,000 40,000
Product line segment margin 532,000 196,000 336,000
Common fixed expenses 100,000
Net operating income 432,000